OpenStudy (anonymous):

how do i use the geometric sequence formula to figure out what the 6th term is (after 7)? 15,13,11,9,7 ?

OpenStudy (anonymous):

(15-n)=13 what is n

OpenStudy (anonymous):

so you take 7-2=5 the 6th term is 5

OpenStudy (anonymous):

yes, but i need to use the geometric sequence formula

OpenStudy (anonymous):

\[a_n=a_1r ^{n-1}\]

OpenStudy (anonymous):

an=15*2^(n-1)

OpenStudy (anonymous):

does this help or not? @andreadesirepen

OpenStudy (anonymous):

n-1 would be 5?

OpenStudy (anonymous):

wouldnt the r be 1.2?

OpenStudy (jtvatsim):

Hmm... sorry to contradict, but this doesn't look like a geometric sequence at all, it looks like an aritmetic sequence.

OpenStudy (jtvatsim):

geometric sequences are the result of multiplying (but there's nothing being constantly multiplied here).

OpenStudy (anonymous):

this is what im being asked: What is the height of each ball on the fifth bounce (i.e., Height 6)? Use the geometric sequence formula

OpenStudy (anonymous):

i dropped the ball from 15ft high, and the first time it bounced 13ft ft then 11 then 9 etc...

OpenStudy (anonymous):

@jtvatsim

OpenStudy (jtvatsim):

hmm.. it just doesn't make sense because your ratio is changing every time, so you can't write a typical geometric sequence...

OpenStudy (jtvatsim):

your ratio starts as 13/15 = 0.867 Then changes to 11/13 = 0.846. Then changes again to 9/11 = 0.818. If the ratio isn't constant, I'm not sure that it's possible to even use a geometric sequence.

OpenStudy (anonymous):

what if i round if off and leave it at just 8?

OpenStudy (anonymous):

*.8

OpenStudy (jtvatsim):

perhaps, that would be an approximation, so I hope an exact answer isn't needed. -.-

OpenStudy (anonymous):

i thought it had to be 15/ 13 though

OpenStudy (jtvatsim):

I'm pretty sure that it has to be 13/15 because your values are getting smaller over time.

OpenStudy (jtvatsim):

If they were growing then 15/13 would make sense.

OpenStudy (anonymous):

oh ok. so i would use this equation, 15*(.8)^5 ?

OpenStudy (jtvatsim):

yes